how can i get standard deviation and sample size from the following data
Data from five paediatric centres were compared in children
who received 0.05 (41 episodes) or 0.1 units/kg/h (52 episodes)
In the low vs. standard dose group, at 6 h following admission,
the fall in blood glucose levels [11.3 (95% confidence interval 8.6 to 13.9) vs. 11.8 (8.4 to 15.2) mmol/L, p = 0.86] and rise in pH [0.13 (0.09 to 0.18) vs.
0.11 (0.07 to 0.15), p = 0.78] were similar
it's not home work it's a research

Answer by rothauserc(4718)

Is p = 0.86 a statistics p-value? If not, what is it?
:
if we assume a standard normal distribution
:
**************************
consider the first two confidence intervals with p = 0.86
:
upper 95% limit = sample mean + (standard error for sample mean(SEM) * 1.96)
:
13.9 = 11.3 + (SEM * 1.96)
:
SEM = (13.9 - 11.3) / 1.96 = 1.326530612
:
sample size = (critical value * SEM / margin of error (ME))^2
:
note that ME for a 95% confidence interval is 0.05
:
sample size = (1.96 * 1.326530612 / 0.05)^2 is approx 2704